OpenStudy (anonymous):

What are 3 solutions on a graph to 2y=5x+11

OpenStudy (anonymous):

well you can find out by changing this equation into the slope intercept form y = 5/2 x + 11/2

OpenStudy (anonymous):

Welcome to Openstudy. Basically what they mean by solutions are points on the graph

OpenStudy (anonymous):

So we plug in any 3 values for x (say -1, 0, and 1) and plug them into the "x" part of our equation and solve for y

OpenStudy (anonymous):

Yeah but I need help doing

OpenStudy (anonymous):

Can you give me 3 ordered pairs that are correct?(: Everything I am doing is always wrong on this and I really need too pass this!

OpenStudy (anonymous):

x= -1 2y=5(-1)+11 2y= -5+11 2y=6 y=3 so we plugged in -1 for x and got a y-value of 3, so one solution is (-1,3) We need to do the same, except plug in 0 and 1 If you don't follow just let me know (:

OpenStudy (anonymous):

Thank you! is 1,8 a solution? Can you give me 1 more ordered pair lol

OpenStudy (anonymous):

Yes it is! You can pick anything you want for x

OpenStudy (anonymous):

I messaged you lol

OpenStudy (anonymous):

Oh yeah, you just have to plug in any number in for x and then solve for y, and then those two numbers are the order pair, where the x goes first and then the y goes second

OpenStudy (anonymous):

The same method you used to get (1,8), except instead of using 1, you can use any number, I find that using that 0 is relatively easy